OpenStudy (anonymous):

f(x) = −3x^2 − 18x − 29 put into a(x-h)^2+k form

OpenStudy (anonymous):

you are supposed to complete the square to find the answer. so far i have -3(x2-6x+9)-29-9

OpenStudy (anonymous):

i have to find the standard equation for a parabola with vertical axis

OpenStudy (anonymous):

a(x-h)² +k =ax²-2axh+ah²+k=-3x²-18x-29 so a=-3 -2ah=-18 ah²+k=-29

OpenStudy (anonymous):

the equality of two polynomes , it's a mathematic propriety ; you will find a , h and k

OpenStudy (anonymous):

? that is not the way my book shows

OpenStudy (anonymous):

factor out the \(-3\) get \[(x)=-3(x^2+6x)-29\] the half of 6 is 3 and write \[f(x)=-3(x+3)^2+k\] find \(k\) by replacing \(x\) by \(-3\) in the original function

OpenStudy (anonymous):

the problem with the method you wrote as a first step, is when you completed the square you thought you added a 9 and therefore subtracted a 9 but in fact you did not add 9 you subtracted 27 (distributive property) so you have to add 27

OpenStudy (anonymous):

a=-3 h=-3 k=-2 by solving systema

OpenStudy (anonymous):

you wrote \[-3(x^2-6x+9)-29-9\] but the +9 is actually \(-27\) since it is being multiplied by \(-3\)

OpenStudy (anonymous):

you should write \[-3(x^2+6x+9)-29+27\]

OpenStudy (anonymous):

-3(x+3)^2-27?

OpenStudy (anonymous):

notice also that when you factor out the \(-3\) is not \(-6x\) but rather \(+6x\)

OpenStudy (anonymous):

there is this method ; of equality of two polynomes

OpenStudy (anonymous):

\(-29+27=-2\)

OpenStudy (anonymous):

\[f(x)=-3(x+3)^2-2\] should do it that is why i think it is easier to evaluate the function at \(-3\) then to try to keep track of the constant

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you can avoid the common mistake of not realizing what you added or subtracted by just starting with \[f(x)=-3(x+3)^2+k\] and since clearly \[f(-3)=k\] you can find \(k\) by finding \(f(-3)\) using the original expression for \(f(x)\)